A rock dropped into a well is heard to splash 2.30 seconds later. How deep is the well to the water's surface?
1
Expert's answer
2017-12-06T13:09:06-0500
s=(gt^2)/2=1/2×9.81 m/s^2 × 2.30 s^2≈11.28 m